OpenStudy (anonymous):

Identify the vertex for the graph of y = −2x2 − 12x + 1.

Nnesha (nnesha):

formula to find x-coordinate of the vertex \[\rm x=\frac{-b}{2a}\] and then substitute x for its value into the equation to find y-coordiante

Nnesha (nnesha):

y=Ax^2+Bx+C quadratic equation where a=leading coefficient b=middle term(coefficient of x term) c=constant

OpenStudy (anonymous):

(−3, 19) (−3, 55) (3, −53) (3, −47)

OpenStudy (anonymous):

those are the options

Nnesha (nnesha):

okay.good now plugin a b values into the formula to find x-coordinate

OpenStudy (anonymous):

Is it C?

Nnesha (nnesha):

how did you get that ? post ur work

OpenStudy (anonymous):

I did it on my notebook its to long, it is correct right?

Nnesha (nnesha):

take a pic post it here it should't be that long what's a value what is b ?

OpenStudy (whpalmer4):

You can check your work pretty easily by evaluating two more points. You think that \((3,-53)\) is the vertex. If it is the vertex, then the points at \(x=3+1=4\) and \(x=3-1=2\) ought to have the same \(y\) value, because the parabola is symmetrical about a line through the vertex. Plug in those two values of \(x\) in the original formula and see if you get identical values for \(y\). If you do, you are correct. If you do not, you are not.

OpenStudy (whpalmer4):

Hint: I strongly suggest you use my test :-)